OpenStudy (anonymous):

Find the constant difference for a hyperbola with foci at F1 (-29, 0) and F2 (29, 0), and point on the hyperbola (21, 0). 42 36 39 45?

OpenStudy (dumbcow):

not sure what you mean by constant difference?

OpenStudy (dumbcow):

the point (21,0) must be a vertex so a = 21 c = 29 b = sqrt(29^2 - 21^2) = 20 hyperbola: \[\frac{x^{2}}{21^{2}} -\frac{y^{2}}{20^{2}} = 1\]

OpenStudy (dumbcow):

the foci are defined such that the difference between distances from any point and each foci is 2a 2a = 42